### v1

@@ III-C2 Transferring Category Semantics Across Domains via Dual-Branch Networks
Subsequently, we constructed two feature matrices: one initialized with the category text features learned from the source domain, denoted as $G^{S}$ with a shape of [C, D], where C is the number of categories and D is the dimension of the category text features. This matrix, containing semantic information learned from the source domain, is frozen during training. The other is a matrix $W_{e}$ initialized with high-confidence sample image features $F^{\prime}$ from the target domain, which has a shape of [C, K, D]. Here, $K$ represents the number of high-confidence samples per category. This matrix is subject to parameter updates during training.
Next, we use a cross-attention mechanism to fuse the target domain image feature matrix $W_{e}$ with the class text feature matrix $G^{S}$ learned from the source domain. During the fusion process, $G^{S}$ and $W_{e}$ are passed through independent linear layers and then used as the queries (Q) and keys/values (K, V) in the cross-attention layer, respectively. This approach allows each class text feature in $G^{S}$ to be enhanced by $K$ high-confidence samples of the corresponding class from $W_{e}$. By encouraging the class text features to attend to the features of different image samples in the target domain, this improves the classification performance of the class text features on target samples.
The attention output after $W_{e}$ and $G^{S}$ pass through the cross-attention layer is computed as follows:
@@
Although the performance of the enhanced source domain class text features $W_{g}^{*}$ improves in the target domain, to further learn target-specific class text features, we introduce a branch that adapts CLIP to the target domain through learnable class-shared soft prompts, as shown in Figure 2. Specifically, we set up a learnable sequence of class-shared tokens $T_{t}=[V_{1}^{T},V_{2}^{T},\ldots,V_{M}^{T}]$, where $V_{m}^{T}\in\mathbb{R}^{D}$ and $M$ is the predefined number of learnable tokens. During prediction, we use this token sequence to replace the original text prompt to obtain the class text features $g_{c}^{T}$. For a target domain sample $x_{i}^{T}$, the classification result of this branch is given by:
@@
where $\delta$ represents the softmax function, $\alpha$ is a predefined hyperparameter between 0 and 1, set to 0.5 by default in this work.
III-D Unsupervised Optimization Strategy Driven by Accurate Classification and Diversity
As described in Section III-C2, the strategy proposed in this paper differs from most Source-Free Unsupervised Domain Adaptation (SF-UDA) methods in that it does not directly fine-tune a model trained in the source domain. Instead, we use class textual features learned from the source domain as a source of semantic information within a cross-domain feature transfer branch to balance semantic information from both the source and target domains during target domain training. This approach allows the model to not only learn classification outcomes for source domain categories but also to acquire semantic features specific to the target domain through a dedicated feature learning branch. However, this strategy might result in a lack of fundamental classification capabilities at the initial stages of target domain training. To address this shortcoming, we introduce an Unsupervised Optimization Strategy Driven by Accurate Classification and Diversity, which aims to maintain the classification abilities learned from the source domain while further enhancing the model’s classification performance in the target domain.
Specifically, we utilize the high-confidence sample set $D_{T}^{\prime}$, filtered as mentioned in Section III-C2, as a reliable collection with pseudo-labels. The classification capability of the source domain model is restored through the computation of cross-entropy loss between these samples and their pseudo-labels. The formula for this computation is as follows:
@@ III-D Unsupervised Optimization Strategy Driven by Accurate Classification and Diversity
Although $\mathcal{L}_{\text{consistency}}$ enhances the diversity of the model’s learned class feature distribution, it fundamentally relies on pseudo-label information, which might bias the model towards categories that are easier to classify. For samples that are inherently more difficult to distinguish, the model may still make classification errors. To address this issue, inspired by previous work [23, 51], we aim to maximize the mutual information between the model’s output class distribution and the input data. This approach encourages the model’s outputs to exhibit both individual certainty and global diversity, thus reducing the reliance on potentially inaccurate pseudo-labels and improving the robustness of the classification across more complex or ambiguous categories.
@@ IV-A1 Datasets
To prove the effectiveness of our proposed method, we refer to the previous method [23, 25] and select four public standard domain adaptation datasets and three publicly available remote sensing domain scene recognition datasets to evaluate our methods, which are Office-31 [31], Office-Home [46], VisDA [27], DomainNet-126  [28], UC Merced Land-Used [52], NWPU-RESISC45 [1] and AID [49], Office-31 and Office-Home are small and medium-sized datasets, VisDA and DomainNet-126 are both large-scale datasets, and the latter two are more challenging for Source-free domain adaptation tasks. UCM Merced Land-Use is a small remote sensing image dataset, while NWPU-RESISC45 and AID are two larger datasets. Overall, we validate the effectiveness of our method on these seven datasets with thirty-seven different transfer tasks.
@@
DomainNet-126 is another large-scale datasets collected from 4 domains, Real(Rl), Sketch(Sk), Clipart(Cl) and Painting(Pn). In the previous source free domain adaptation tasks, they usually evaluate the methods on 7 transfer tasks due to the severe noisy labels in the dataset [32]. Now, we can use the clean version of the dataset and evaluate our method on 12 transfer tasks.
UC Merced Land-Use, abbreviated as UCM hereafter, is a remote sensing dataset released in 2010 by the Computer Vision Laboratory at UC Merced, Moosehead. It is widely used for remote scene recognition to categorize land-use scenes in urban areas. The dataset contains 21 categories, each with 100 remote sensing images of size 256x256 pixels, amounting to a total of $2{,}100$ images.
NWPU-RESISC45, referred to as NWPU in this document, is a benchmark for remote sensing scene recognition, developed by Northwestern Polytechnical University (NWPU) in 2017. It consists of 45 scene categories, each containing 700 images, for a total of $31{,}500$ images, all with a resolution of 256x256 pixels.
@@
Ours no 73.9 94.1 91.6 85.7 94.3 91.7 84.9 73.7 91.9 85.5 72.7 94.0 86.2
TABLE I: Performance comparison on OfficeHome.
Method SF Source $f_{v}$ Cl→Pn Cl→Rl Cl→Sk Pn→Cl Pn→Rl Pn→Sk Rl→Cl Rl→Pn Rl→Sk Sk→Cl Sk→Pn Sk→Rl Avg
@@
Ours no 84.8 92.4 81.9 87.6 93.0 82.1 87.5 84.9 81.6 87.9 84.8 91.9 86.7
TABLE II: Performance comparison on Mini-DomainNet.
Method SF Source $f_{v}$ A→D A→W D→A D→W W→A W→D Avg.
@@
Ours no 94.8 95.8 85.1 95.8 85.6 96.6 92.3
TABLE III: Performance comparison on Office-31.
Method SF Source $f_{v}$ plane bcycl bus car horse knife mcycle person plant sktbrd train truck Avg.
@@
Ours no 97.9 94.0 88.1 70.3 98.4 95.0 93.4 82.5 91.1 95.9 92.9 78.9 89.9
TABLE IV: Performance comparison on VisDA-2017.
@@ IV-A2 Implementation Details
We utilized a pre-trained CLIP model as the backbone network and compared two image encoder architectures: ResNet-50[13] and ViT-B/16[5]. In experiments conducted on the VisDA-2017 dataset, we employed ResNet-101 to ensure a fair comparison with other methods. The optimizer used was Adam, with a momentum of 0.9 and a weight decay set to $5\times 10^{-4}$. The initial learning rate was set at 0.001. During the adaptive training phase, we implemented a cosine annealing learning rate scheduler for dynamic adjustment. In our experiments, the batch size was set to 32, with $\alpha$ equal to 0.5 and $M$ set to 16. All experiments were designed and implemented using the PyTorch platform.
@@ IV-B1 Competitors
In this study, we explore a relatively under-investigated area: Source-Free Unsupervised Domain Adaptation (SF-UDA) under conditions of extremely limited source domain samples. Given the uniqueness of this novel setting, we find it challenging to identify existing studies with a completely analogous setup for comparison. Therefore, to validate the effectiveness of our method, we chose to compare it against SF-UDA methods that train with a large number of source domain samples and UDA methods that have direct access to substantial source domain samples. This comparison not only demonstrates the adaptability of our approach under the dual challenges of no direct access to source domain data and limited source sample availability but also underscores its innovation and practicality. By doing so, we aim to illustrate that even with minimal source data, our approach can still perform comparably to methods that benefit from richer source environments, highlighting its potential for scenarios where traditional UDA assumptions do not hold.
@@
  • Twelve SFDA methods, which, although unable to directly access source domain data, still operate with a relatively abundant number of source domain samples.
@@ IV-B2 Experimental Results
As demonstrated in Table I and II, our method achieve the best performance of 78.1% and 77.3% on the standard unsupervised domain adaptation datasets e Office-Home and Mini-DomainNet, respectively, utilizing ResNet-50 as the image backbone network. These results surpassed previous Source-Free Unsupervised Domain Adaptation (SF-UDA) methods and standard UDA methods. This outcome not only reflects our model’s capability to maintain high performance in the target domain despite the dual challenge of extremely limited source domain samples but also highlights its superiority.
In scenarios where the image backbone network is Vit-B/16, although our method does not directly access source domain samples, it outperforms most direct source domain-accessing CLIP-based UDA methods and performs comparably to the currently best-performing PAD-CLIP method.
Table III and IV further showcase that our method reaches performance levels on the Office31 and VisDA-2017 datasets comparable to those of previous SF-UDA methods that utilize a large number of source domain samples. This further validates the efficiency of our approach in utilizing data.
Specifically in professional applications, as illustrated in Table V, our method achieves performance similar to DATSNET, a UDA method specifically designed for remote sensing images, in domain adaptation tasks targeting remote sensing imagery. This achievement is particularly noteworthy given that we use only a limited number of source domain samples. In the field of remote sensing, the gap between source and target domains is typically larger than in standard DA datasets, making source samples crucial for training models that generalize well to the target domain. Moreover, compared to other UDA methods that utilize CLIP and source domain data such as DAPL and AD-CLIP, our method exhibits a clear performance advantage, exceeding AD-CLIP by 3.9%.
@@
Ours no 92.9 98.6 93.8 92.3 97.3 91.5 94.4
TABLE V: Accuracy (%) for cross-domain remote scene recognition (UCM, AID, and NWPU).
Source training Target Prompt Fusion Module Trainable cache key Avg
@@
IV-C Alblations Studies
The Effectiveness of Different Components in the Model Architecture. In Table VI, we analyze the effectiveness of different components in the model using the e Office-Home dataset, comparing the average accuracy of four distinct model configurations across the four domains. Initially, when trained solely on source domain samples, the model achieves an accuracy of 82.5%. This indicates that learning class labels from a limited number of source domain samples retain substantial generalization information relevant to the target domain, preventing overfitting to the source domain feature distribution. When the model undergoes unsupervised training using only soft prompts learned from the target domain, accuracy improves to 84.3%, confirming the effectiveness of our proposed unsupervised optimization strategy in maintaining high performance even with exclusive target domain training. Furthermore, incorporating a fusion module to integrate semantic category information from the source domain into the target domain training raises model performance to 85.7%, demonstrating that category semantic information learned from a small number of source domain samples effectively supports training in the target domain. Finally, setting the feature matrix of high-confidence target domain samples, $W_{e}$, as a learnable parameter further enhances model performance to 86.2%. These results collectively highlight the contributions of each component to enhancing the model’s generalization ability.
@@ IV-C Alblations Studies
The Effectiveness of Different Loss Terms in Unsupervised Optimization Strategies. In Table VII, we analyze the effectiveness of the three loss components in our proposed unsupervised optimization strategy across the e Office-Home, VisDA-2017, and Mini-DomainNet datasets. By comparing the impact of different combinations of loss components on the accuracy of these datasets, we observe that the cross-entropy loss for high-confidence samples, $\mathcal{L}_{CE}^{T}$, and the mutual information maximization loss, $\mathcal{L}_{IM}$, significantly enhance the model’s performance. This effect may stem from the fact that both $\mathcal{L}_{CE}^{T}$ and $\mathcal{L}_{IM}$ promote individual certainty in model predictions and global diversity. However, since $\mathcal{L}_{IM}$ computes the loss over all samples, while $\mathcal{L}_{CE}^{T}$ only considers a small number of highly reliable samples, the effect of $\mathcal{L}_{IM}$ is relatively more pronounced. Furthermore, the multi-view consistency loss, $\mathcal{L}_{\text{consistency}}$, improves model performance regardless of its combination with other losses; when all three loss components are used together, the model achieves optimal performance, indicating that each component contributes effectively.
The Effectiveness of Logits Fusion in Dual-Branch Networks. As shown in Figure 3, we randomly selected the migration task Product → Real World from the e Office-Home dataset and visualized the accuracy changes of the cross-domain transfer branch and the target-specific feature learning branch over the training epochs. The results indicate that the accuracy of both branches gradually improves as training progresses. Furthermore, the accuracy of the merged logits for both branches consistently exceeds that of any single branch in each epoch. This demonstrates the effectiveness of our proposed dual-branch logits merging strategy in enhancing model performance.
@@
Impact of Source Domain Sample Size. As illustrated in Table VIII, we analyzed the impact of the number of samples in the source domain on the performance of the transferred model in the target domain across three remote sensing datasets. It is evident that when the number of samples per class in the source domain is fewer than eight, increasing the sample size significantly enhances the model’s performance. Specifically, the performance improved by 6.82% when the sample count increased from one to eight per class. This improvement can be attributed to the more comprehensive semantic information about the class contained within the samples at this stage. However, when the sample count reaches sixteen per class, the model’s performance does not improve and slightly decreases compared to having eight samples per class. This decline is likely due to the semantic features learned being overly fitted to the source domain, and during the training of the source domain samples, the amount of learnable parameters was limited. Thus, when the number of source domain samples exceeds eight, the useful semantic information for the target domain has likely reached saturation.

### v2

@@ III-D Cross-Modal Dual-Branch Networks
Then, we constructed two feature matrices: one initialized with the category text features learned from the source domain, denoted as $G^{S}$ with a shape of [C, D], where C is the number of categories and D is the dimension of the category text features. This matrix, containing semantic information learned from the source domain, is frozen during training. The other is a matrix $W_{e}$ initialized with high-confidence sample image features $F^{\prime}$ from the target domain, which has a shape of [C, K, D]. Here, $K$ represents the number of high-confidence samples per category. This matrix is modified by parameter updates during training.
Next, we use a cross-attention module with residual connections to fuse the target domain image feature matrix $W_{e}$ with the class text feature matrix $G^{S}$ learned from the source domain. During the fusion process, $G^{S}$ and $W_{e}$ are passed through independent linear layers and then used as the queries (Q) and keys/values (K, V) in the cross-attention layer, respectively. This approach allows each class text feature in $G^{S}$ to be enhanced by $K$ high-confidence samples of the corresponding class from $W_{e}$. By encouraging the class text features to attend to the features of different image samples in the target domain, this improves the classification performance of the class text features on target samples.
The attention output after $W_{e}$ and $G^{S}$ passes through the cross-attention layer is computed as follows:
@@
Although the performance of the enhanced source domain class text features $W_{g}^{*}$ improves in the target domain, to further learn target-specific class text features, we introduce a branch that adapts CLIP to the target domain through learnable class-shared soft prompts, as shown in Figure 2. Specifically, we set up a learnable sequence of class-shared tokens $T_{t}=[V_{1}^{T},V_{2}^{T},\ldots,V_{M}^{T}]$, where $V_{m}^{T}\in\mathbb{R}^{D}$ and $M$ are the predefined numbers of learnable tokens. During prediction, we use this token sequence to replace the original text prompt to obtain the class text features $g_{c}^{T}$. For a target domain sample $x_{i}^{T}$, the classification result of this branch is given by:
@@
where $\delta$ represents the softmax function, $\alpha$ is a predefined hyper-parameter between 0 and 1, set to 0.5 by default in this work.
III-E Unsupervised Optimization Strategy Driven by Accurate Classification and Diversity
As described in Section III-D, the strategy proposed in this paper differs from most Source-Free Unsupervised Domain Adaptation (SF-UDA) methods in that it does not directly fine-tune a model trained in the source domain. Instead, we use class textual features learned from the source domain as a source of semantic information within a cross-domain feature transfer branch to balance semantic information from both the source and target domains during target domain training. This method enables the model to effectively transfer semantic information from the source domain while minimizing the noise and domain shift effects present in the feature distribution learned from a limited number of source domain samples. However, this strategy might result in a lack of fundamental classification capabilities at the initial stages of target domain training. To address this shortcoming, we introduce an Unsupervised Optimization Strategy Driven by Accurate Classification and Diversity, which aims to maintain the classification abilities learned from the source domain while further enhancing the model’s classification performance in the target domain.
Specifically, we utilize the high-confidence sample set $D_{T}^{\prime}$, filtered as mentioned in Section III-D, as a reliable collection with pseudo-labels. The classification capability of the source domain model is restored through the computation of cross-entropy loss between these samples and their pseudo-labels. The formula for this computation is as follows:
@@ III-E Unsupervised Optimization Strategy Driven by Accurate Classification and Diversity
Although $\mathcal{L}_{\text{consistency}}$ enhances the diversity of the model’s learned class feature distribution, it fundamentally relies on pseudo-label information, which might bias the model towards categories that are easier to classify. For samples that are inherently more difficult to distinguish, the model may still make classification errors. To address this issue, and inspired by previous work [24, 52], we aim to maximize the mutual information between the model’s output class distribution and the input data. This approach encourages the model’s outputs to exhibit both individual certainty and global diversity, thus reducing the reliance on potentially inaccurate pseudo-labels and improving the robustness of the classification across more complex or ambiguous categories.
@@ IV-A1 Datasets
To prove the effectiveness of our proposed method, we refer to the previous method [24, 26] and select four public standard domain adaptation datasets and three publicly available remote sensing domain scene recognition datasets to evaluate our methods, which are Office-31 [32], Office-Home [47], VisDA [28], Mini-DomainNet [29], UC Merced Land-Use [53], NWPU-RESISC45 [2] and AID [50]. Office-31 and Office-Home are small and medium-sized datasets, VisDA and Mini-DomainNet are both large-scale datasets, and the latter two are more challenging for Source-free domain adaptation tasks. UC Merced Land-Use is a small remote sensing image dataset, while NWPU-RESISC45 and AID are two larger datasets. Overall, we validate the effectiveness of our method on these 7 datasets with 31 different transfer tasks.
@@
Mini-DomainNet is another large-scale dataset collected from 4 domains, Real(Rl), Sketch(Sk), Clipart(Cl) and Painting(Pn). In the previous source-free domain adaptation tasks, they usually evaluate the methods on 7 transfer tasks due to the severe noisy labels in the dataset [33]. Now, we can use the clean version of the dataset and evaluate our method on 12 transfer tasks.
UC Merced Land-Use, abbreviated as UCM hereafter, is a remote sensing dataset released in 2010 by the Computer Vision Laboratory at UC Merced. It is widely used for remote scene recognition to categorize land-use scenes in urban areas. The dataset contains 21 categories, each with 100 remote sensing images of size 256x256 pixels, amounting to a total of $2{,}100$ images.
NWPU-RESISC45, referred to as NWPU in this document, is a benchmark for remote sensing scene recognition, developed by Northwestern Polytechnical University (NWPU) in 2017. It consists of 45 scene categories, each containing 700 images, for a total of $31{,}500$ images, all with a resolution of $256\times 256$ pixels.
@@
Ours 8-shots 73.9 94.1 91.6 85.7 94.3 91.7 84.9 73.7 91.9 85.5 72.7 94.0 86.2
TABLE I: Accuracy (%) of Different Settings and Domain Adaptation Methods on the Office-Home Dataset [47]. SF indicates the accessibility of the source domain, Source refers to the number of samples in the source domain, and $E_{img}$ denotes the image encoder used. The highest accuracy for each encoder is highlighted in bold, while the second highest is underlined for clarity.
Method SF Source $E_{img}$ Cl→Pn Cl→Rl Cl→Sk Pn→Cl Pn→Rl Pn→Sk Rl→Cl Rl→Pn Rl→Sk Sk→Cl Sk→Pn Sk→Rl Avg
@@
Ours 8-shots 84.8 92.4 81.9 87.6 93.0 82.1 87.5 84.9 81.6 87.9 84.8 91.9 86.7
TABLE II: Accuracy (%) of Different Settings and Domain Adaptation Methods on the Mini-DomainNet[29] dataset.
Method SF Source $E_{img}$ A→D A→W D→A D→W W→A W→D Avg.
@@
Ours 8-shots 94.8 95.8 85.1 95.8 85.6 96.6 92.3
TABLE III: Accuracy (%) of Different Settings and Domain Adaptation Methods on the Office-31[32] Dataset.
Method SF Source $E_{img}$ plane bcycl bus car horse knife mcycle person plant sktbrd train truck Avg.
@@
Ours 8-shots 97.9 94.0 88.1 70.3 98.4 95.0 93.4 82.5 91.1 95.9 92.9 78.9 89.9
TABLE IV: Accuracy (%) of Different Settings and Domain Adaptation Methods on the VisDA-2017[28] Dataset.
@@ IV-A2 Implementation Details
We utilized a pre-trained CLIP model as the backbone network and compared two image encoder architectures: ResNet-50[14] and ViT-B/16[6]. In experiments conducted on the VisDA-2017 dataset, we employed ResNet-101 to ensure a fair comparison with other methods. The optimizer used was Adam, with a momentum of 0.9 and a weight decay set to $5\times 10^{-4}$. The initial learning rate was set at 0.001. During the adaptive training phase, we implemented a cosine annealing learning rate scheduler for dynamic adjustment. In our experiments, the batch size was set to 32, with $\alpha$ set to 0.5, $K$ set to 8, and $M$ set to 16. During the experiment, we randomly selected 8 samples from each category in the source domain and calculated the average of the results obtained under three different random seeds as the final result. All experiments were designed and implemented using the PyTorch platform.
@@ IV-B1 Competitors
In this study, we explore a relatively under-investigated area: Source-Free Unsupervised Domain Adaptation (SF-UDA) under conditions of extremely limited source domain samples. Given the uniqueness of this novel setting, we find it challenging to identify existing studies with a completely analogous setup for comparison. Therefore, to validate the effectiveness of our method, we chose to compare it against SF-UDA methods that train with a large number of source domain samples and UDA methods that have direct access to substantial source domain samples. This comparison not only demonstrates the adaptability of our approach under the dual challenges of no direct access to source domain data and limited source sample availability but also highlights its innovation and practicality. By doing so, we aim to illustrate that even with minimal source data, our approach can still perform comparably to methods that benefit from richer source environments, highlighting its potential for scenarios where traditional UDA assumptions do not hold.
@@
  • Twelve SF-UDA methods, which, although unable to directly access source domain data, can still utilize a relatively rich number of source domain samples.
@@ IV-B2 Experimental Results
First, as shown in Table I and Table II, our method achieved the best performance on the standard unsupervised domain adaptation datasets Office-Home and Mini-DomainNet, with 78.1% and 77.3% respectively, using ResNet-50 as the backbone network. These results outperform previous source-free unsupervised domain adaptation (SF-UDA) methods as well as standard UDA approaches, clearly demonstrating that our method can achieve top performance in the target domain even under strict source data access constraints, showcasing its efficiency and robustness. Additionally, while our method did not surpass all approaches when using ViT-B/16 as the backbone, it achieved results nearly identical to the best-performing UDA method, PAD-CLIP, further validating its competitiveness.
Second, as shown in Table III, on the Office-31 dataset, our method, using ViT-B/16 as the image encoder, achieved performance comparable to SF-UDA methods based on large source samples, only 0.7% lower than DSiT-B. Although ResNet-50 showed relatively weaker pretraining performance, our method still demonstrated its effectiveness under limited source sample conditions, proving its adaptability on datasets with smaller domain gaps.
Finally, as shown in Table IV and Table V, when using ViT-B/16 as the backbone network, our method surpassed previous SF-UDA methods on the VisDA-2017 dataset, trailing the best UDA method, PAD-CLIP, by only 1.0%. Additionally, it achieved performance comparable to DATSNET, a UDA method specifically designed for remote sensing images, in domain adaptation tasks involving remote sensing imagery. These results not only confirm the effectiveness of our method but also highlight its significant performance advantage in professional applications with limited source samples, particularly by outperforming AD-CLIP by 3.3%, demonstrating strong competitiveness.
These experimental results collectively validate that our proposed method can maintain strong performance across various tasks and settings, even under stringent conditions with limited source data, showcasing its robustness and generalizability.
@@
Ours no 92.3 98.8 92.0 91.6 96.3 91.5 93.8
TABLE V: Accuracy (%) for cross-domain remote scene recognition on the UCM[53], AID[50], and NWPU[2] datasets.
Source training Target Prompt Fusion dual branches Trainable $W_{e}$ Avg
@@
IV-C Ablations Studies
The Effectiveness of Different Components in the Model Architecture. In Table VI, we analyze the effectiveness of different components in the model using the Office-Home dataset, comparing the average accuracy of four distinct model configurations across the four domains. Initially, when trained solely on source domain samples, the model achieves an accuracy of 82.5%. This indicates that prompt learning from a limited number of source domain samples retains substantial generalization information relevant to the target domain. When the model undergoes unsupervised training using only soft prompts learned from the target domain, accuracy improves to 84.3%, confirming the effectiveness of our proposed unsupervised optimization strategy in maintaining high performance even with exclusive target domain training. Furthermore, incorporating a fusion module to integrate semantic category information from the source domain into the target domain training raises model performance to 85.7%, demonstrating that category semantic information learned from a small number of source domain samples effectively supports training in the target domain. Finally, setting the feature matrix of high-confidence target domain samples, $W_{e}$, as a learnable parameter further enhances model performance to 86.2%. These results collectively highlight the contributions of each component to enhancing the model’s generalization ability.
@@ IV-C Ablations Studies
The Effectiveness of Different Loss Terms in Unsupervised Optimization Strategies. In Table VII, we analyze the effectiveness of the three loss components in our proposed unsupervised optimization strategy across the Office-Home, VisDA-2017, and Mini-DomainNet datasets. By comparing the impact of different combinations of loss components on the accuracy of these datasets, we observe that the cross-entropy loss for high-confidence samples, $\mathcal{L}_{CE}^{T}$, and the mutual information maximization loss, $\mathcal{L}_{IM}$, significantly enhance the model’s performance. This effect may stem from the fact that both $\mathcal{L}_{CE}^{T}$ and $\mathcal{L}_{IM}$ promote individual certainty in model predictions and global diversity. However, since $\mathcal{L}_{IM}$ computes the loss over all samples, while $\mathcal{L}_{CE}^{T}$ only considers a small number of highly reliable samples, the effect of $\mathcal{L}_{IM}$ is relatively more pronounced. Furthermore, the multi-view consistency loss, $\mathcal{L}_{\text{consistency}}$, improves model performance regardless of its combination with other losses; when all three loss components are used together, the model achieves optimal performance, indicating that each component contributes effectively.
The Effectiveness of Logits Fusion in Dual-Branch Networks. As shown in Figure 3, we randomly selected the migration task Product → Real World from the Office-Home dataset and visualized the accuracy changes of the cross-domain transfer branch and the target-specific feature learning branch over the training epochs. The results indicate that the accuracy of both branches gradually improves as training progresses. Furthermore, the accuracy of the merged logits for both branches consistently exceeds that of any single branch in each epoch. This demonstrates the effectiveness of our proposed dual-branch logits merging strategy in enhancing model performance.
@@
Impact of Source Domain Sample Size. As illustrated in Table VIII, we analyzed the impact of the number of samples in the source domain on the performance of the transferred model in the target domain across three remote sensing datasets. It is evident that when the number of samples per class in the source domain is fewer than eight, increasing the sample size significantly enhances the model’s performance. Specifically, the performance improved by 6.35% when the sample count increased from one to eight per class. This improvement can be attributed to the more comprehensive semantic information about the class contained within the samples at this stage. However, when the sample count reaches sixteen per class, the model’s performance does not improve and slightly decreases compared to having eight samples per class. This decline is likely due to the semantic features learned being overfitted to the source domain, and during the training of the source domain samples, the amount of learnable parameters was limited. Thus, when the number of source domain samples exceeds eight, the useful semantic information for the target domain has likely reached saturation.